OpenStudy (lgbasallote):

\[\huge a^2 (dy - dx) = x^2 dy + y^2 dx; \text{a constant}\] ??

OpenStudy (ash2326):

\[ a^2 (dy - dx) = x^2 dy + y^2 dx\] rewrite it, you'd get \[dy(a^2-x^2)=dx(a^2+y^2)\] \[\frac{dy}{y^2+a^2}=\frac{dx}{a^2-x^2}\] Now integrate both sides, do you need help with it?

OpenStudy (ash2326):

@lgbasallote

OpenStudy (lgbasallote):

i see..clever..first one is arctan right? second is arcsin?